Find an equation for the line with the given properties. Express the equation in slope-intercept form.
slope=9; containing the point (6, -3)
What is the equation of the line?
y is ?

Answer by ankor@dixie-net.com(22740)

slope=9; containing the point (6, -3)
What is the equation of the line?
:
First use the point/slope form y - y1 = m(x - x1)
m = 9, x1 = 3, y1 = -3
y - (-3) = 9(x - 6)
y + 3 = 9x - 54
y = 9x - 54 - 3
y = 9x - 57, the slope/intercept form